OpenStudy (anonymous):

7. The isotope cesium-137, which has a half-life of 30 years, is a product of nuclear power plants. How long will it take for this isotope to decay to about one-half of its original amount? (a) 0 years (b) 15 years (c) 30 years (d) 60 years @Abhisar @UnkleRhaukus

OpenStudy (anonymous):

So with the formula is B?

OpenStudy (unklerhaukus):

The half-life of a substance is the time it takes for half of it to decay

OpenStudy (abhisar):

\(\huge\frac{1}{2}\) = \(\huge\frac{1}{2^n}\)

OpenStudy (abhisar):

n represents the number of times half time has elapsed.

OpenStudy (anonymous):

so the answer is 15?

OpenStudy (abhisar):

So by solving the above equation u get n= ??

OpenStudy (anonymous):

is n 30 or 137? because if it is 137 then the answer is 68.5

OpenStudy (anonymous):

but if n is 30 the answer is 15

OpenStudy (abhisar):

n = 1

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

1/2^1?

OpenStudy (abhisar):

1/2^1 = 1/2

OpenStudy (abhisar):

So, n=1

OpenStudy (anonymous):

\[\frac{ 1 }{ 2^1 }\]

OpenStudy (abhisar):

this means half life has elapsed once

OpenStudy (abhisar):

Therefore 30 years have elapsed.

OpenStudy (abhisar):

getting my points ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so the answer is C?

OpenStudy (anonymous):

so the answer is in the question

OpenStudy (anonymous):

i don't get how you get the half life of 30 years

OpenStudy (abhisar):

ok wait..brb

OpenStudy (anonymous):

brb

OpenStudy (unklerhaukus):

The question states the isotope has a half-life of 30 years

OpenStudy (abhisar):

The question says that half life of isotope is 30 years

OpenStudy (abhisar):

Now, let the number of times half life (30 years) has elapsed be n. Then total time elapsed will be 30n

OpenStudy (abhisar):

Formula for finding the fraction of isotope left is \(\huge\frac{1}{2^n}\)

OpenStudy (anonymous):

ok

OpenStudy (abhisar):

Question says that \(\huge\frac{1}{2}\) is left.

OpenStudy (anonymous):

which would make the answer 68.5 or 15 years

OpenStudy (abhisar):

This implies that \(\huge\frac{1}{2^n}\)=\(\huge\frac{1}{2}\)

OpenStudy (abhisar):

So n=1

OpenStudy (abhisar):

Therefore total time elapsed = 30*1=30 years

OpenStudy (abhisar):

got it ?

OpenStudy (anonymous):

why is in N=1

OpenStudy (abhisar):

Solve the equation \(\color{blue}{\text{Originally Posted by}}\) @Abhisar This implies that \(\huge\frac{1}{2^n}\)=\(\huge\frac{1}{2}\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

i don't get why or how N=1

OpenStudy (abhisar):

u know how to solve equations ?

OpenStudy (abhisar):

ok leave all

OpenStudy (abhisar):

Lets do it conceptually

OpenStudy (anonymous):

ok

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @UnkleRhaukus The half-life of a substance is the time it takes for half of it to decay \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

Now the question says that only 1/2 is left i.e half is left and half is decayed

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):